# Generative Retrieval with Large Language Models

### v2

@@ Appendix H Experimental Results of Selecting Different Numbers of First Stage Documents
In Figure 7, we conduct experiments to compare the effect of selecting different numbers of first-stage documents (denoted as $k$) in LLM2GR when implemented based on Llama-2-13b. We observe that the impact of $k$ on the final performance is not significant, as the necessary effective reference passages are usually contained within the first few documents. The suboptimal performance observed on the HotpotQA dataset when $k=1$ can be attributed to the dataset requiring two documents to calculate R-Precision.
Appendix I Experimental Results with Different Beam Search Sizes
@@
Figures 8 and 9 show the impact of setting different beam sizes in the first and second stages, respectively, in LLM2GR when implemented based on Llama-2-13b. For the first stage of recalling title identifiers, a larger beam size can achieve better page-level results, thereby slightly improving the effectiveness of the second stage of passage recall. However, in the second stage of fine-grained passage recall, the improvement brought by a larger beam size is not significant and may even lead to a slight decline, possibly due to the introduction of additional noise by a larger beam size.
Appendix J Case Study
@@
In Tables 7 to 12, we present reference cases obtained using the Gold Standard, BM25, and the LLM2GR framework with Llama-2-13b on the NQ, TriviaQA, and HotpotQA datasets. By generating passages more aligned with the question, LLM2GR achieves results that contain the answer in Tables 7, 9, and 11. Table 8 showcases a biology question; although the passage recalled and located by LLM2GR does not contain the annotated answer, it provides a more detailed description of the location and process of pancreatic enzyme cleavage of peptide bonds. However, Tables 10 and 12 show instances where LLM2GR’s recall failed. This is because merely generating a relevant prefix sometimes cannot ensure that the subsequent part will definitely contain the answer, leading to passages that are only broadly related. Ensuring the flexibility of recall and location while considering more subsequent information still requires further exploration. Nevertheless, we can also note that the references obtained by LLM recall are more natural and easier to read compared to those with predefined segmented beginnings. Finally, compared to the NQ dataset, questions in the TriviaQA and HotpotQA datasets are longer, allowing for more vocabulary overlap between questions and passages, which enhances the effectiveness of BM25 sparse retrieval.
@@
Gold Standard: Part II (s.10-21) deals with the marriageable age and the marriage of minors. In the original 1961 Act, marriageable age was set at 16 for females and 18 for males. However, under section 12 of the original 1961 Act a female 14 or 15 years or a male 16 or 17 years could apply to the court for permission to marry. The marriageable age was equalised in 1991 by the “Sex Discrimination Amendment Act 1991", which raised the marriageable age of females to 18.
BM25: marriage. Before 2004, there was no definition of marriage in the 1961 Act, and instead the common law definition used in the English case “Hyde v Hyde" (1866) was considered supreme. Though s.46(1) of the Act required celebrants to explain the legal nature of marriage in Australia to a couple as “the union of a man and a woman to the exclusion of all others, voluntarily entered into for life", these words were descriptive or explanatory, rather than outlining what constituted a legally valid marriage in
LLM2GR: Marriage in Australia is regulated by the federal “Marriage Act 1961" (Cth) , which applies uniformly throughout Australia (including its external territories) to the exclusion of all state laws on the subject. Australian law recognises only monogamous marriages, being marriages of two people, including same-sex marriages, and does not recognise any other forms of union, including traditional Aboriginal marriages, polygamous marriages or concubinage. The marriage age for marriage in Australia is 18 years, but in “unusual and exceptional circumstances" a person aged 16 or 17 can marry with parental consent
Table 7: In the NQ dataset, the Gold Standard, BM25, and LLM2GR reference passage for the query “what is the legal age for marriage in Australia" are provided. The parts containing the answer are highlighted with a grey background.
@@
Gold Standard: “In vivo", chymotrypsin is a proteolytic enzyme (serine protease) acting in the digestive systems of many organisms. It facilitates the cleavage of peptide bonds by a hydrolysis reaction, which despite being thermodynamically favorable, occurs extremely slowly in the absence of a catalyst. The main substrates of chymotrypsin are peptide bonds in which the amino acid N-terminal to the bond is a tryptophan, tyrosine, phenylalanine, or leucine. Like many proteases, chymotrypsin also hydrolyses amide bonds “in vitro", a virtue that enabled the use of substrate analogs such as N-acetyl-L-phenylalanine p-nitrophenyl amide for enzyme assays.
BM25: 149, producing $\alpha$-chymotrypsin (which is more active and stable than $\pi$-chymotrypsin). The resulting molecule is a three-polypeptide molecule interconnected via disulfide bonds. “In vivo", chymotrypsin is a proteolytic enzyme (serine protease) acting in the digestive systems of many organisms. It facilitates the cleavage of peptide bonds by a hydrolysis reaction, which despite being thermodynamically favorable, occurs extremely slowly in the absence of a catalyst. The main substrates of chymotrypsin are peptide bonds in which the amino acid N-terminal
LLM2GR: Chymotrypsin cleaves peptide bonds by attacking the unreactive carbonyl group with a powerful nucleophile, the serine 195 residue located in the active site of the enzyme, which briefly becomes covalently bonded to the substrate, forming an enzyme-substrate intermediate. Along with histidine 57 and aspartic acid 102, this serine residue constitutes the catalytic triad of the active site. These findings rely on inhibition assays and the study of the kinetics of cleavage of the aforementioned substrate, exploiting the fact that the en
Table 8: In the NQ dataset, the Gold Standard, BM25, and LLM2GR reference passage for the query “where does cleavage of the peptide bond by chymotrypsin occur" are provided. The parts containing the answer are highlighted with a grey background.
@@
Gold Standard: Only 193 of the estimated 643 people on board survived, and the soldiers’ chivalry gave rise to the unofficial “women and children first" protocol when abandoning ship, while the “Birkenhead drill" of Rudyard Kipling’s poem came to describe courage in face of hopeless circumstances.
BM25: HMS “Birkenhead, also referred to as HM Troopship “Birkenhead or Steam Frigate “Birkenhead", was one of the first iron-hulled ships built for the Royal Navy. She was designed as a steam frigate, but was converted to a troopship before being commissioned. She was wrecked on 26 February 1852, while transporting troops to Algoa Bay at Danger Point near Gansbaai, 87 miles (140 kilometres) from Cape Town in the Cape Colony. There were not enough serviceable lifeboats for all the passengers, and the soldiers
LLM2GR: The sinking of the “Birkenhead" is one of the earliest maritime disaster evacuations during which the concept of “women and children first" is known to have been applied.“Women and children first" subsequently became standard procedure in relation to the evacuation of sinking ships, in fiction and in life. The term ““Birkenhead" drill" became defined as courageous behaviour in hopeless circumstances and appeared in Rudyard Kipling’s 1893 tribute to the Royal Marines, “Soldier an’ Sailor Too": To take your chance in the thick of a rush, with firing all about, Is nothing so bad when you’ve
Table 9: In the TriviaQA dataset, the Gold Standard, BM25, and LLM2GR reference passage for the query “Which expression is associated with the sinking of the HMS Birkenhead at Gansbaai near Cape Town, South Africa, in Febuary 1852?" are provided. The parts containing the answer are highlighted with a grey background.
@@
Gold Standard: The first series of I’m a Celebrity…Get Me Out of Here! was broadcast on ITV from 25 August to 8 September 2002. Ant & Dec presented the main show on ITV, whilst Louise Loughman hosted the spin-off show “I’m a Celebrity…Get Me Out of Here! NOW!" on ITV2. The winner of this series was radio DJ Tony Blackburn.
BM25: The first series of I’m a Celebrity…Get Me Out of Here! was broadcast on ITV from 25 August to 8 September 2002. Ant & Dec presented the main show on ITV, whilst Louise Loughman hosted the spin-off show “I’m a Celebrity…Get Me Out of Here! NOW!" on ITV2. The winner of this series was radio DJ Tony Blackburn. The show began with 8 celebrity contestants. The contestants take part in daily trials to earn food All ratings are taken from the UK
LLM2GR: I’m a Celebrity…Get Me Out of Here! is a British reality TV series in which a number of celebrities live together in a jungle environment for a number of weeks, competing to be crowned “King" or “Queen of the Jungle". The show was originally created in the United Kingdom by the factual programmes department of ITV’s then London franchise, London Weekend Television and developed by a team including James Allen, Natalka Znak, Brent Baker and Stewart Morris. The first episode aired on 25 August 2002 . It is now produced by ITV Studios and has been licensed globally to countries including the United States
Table 10: In the TriviaQA dataset, the Gold Standard, BM25, and LLM2GR reference passage for the query “Who was the first winner of ‘I’m A Celebrity, Get Me Out Of Here’?" are provided. The parts containing the answer are highlighted with a grey background.
@@
BM25: S is a South Korean project group consisting of three members: Kangta, Lee Ji-hoon and Shin Hye-sung. The group debuted in 2003, under the SM Entertainment label. After 11 years, they released and promoted another mini-album in 2014. In 2003, S released their first album “Fr.In.Cl", which stands for “Friends in Classic". In 2014, after more than a decade in hiatus, the group released their second mini-album “Autumn
LLM2GR: 2014 S/S is the debut album of South Korean group WINNER. It was released on August 12, 2014 by the group’s record label, YG Entertainment. The members were credited for writing the lyrics and composing the majority of the album’s songs. The members produced the majority of the music for this album themselves, with the help of other producers such as Choice 37, B.I., Airplay, and others. The album was highlighted for incorporating elements generally absent from K-pop releases, including hints of acoustic
Table 11: In the HotpotQA dataset, the Gold Standard, BM25, and LLM2GR reference passage for the query “2014 S/S is the debut album of a South Korean boy group that was formed by who?" are provided. The parts containing the answer are highlighted with a grey background.
@@
Gold Standard: Charles Fraser-Smith (26 January 1904 – 9 November 1992) was an author and one-time missionary who is widely credited as being the inspiration for Ian Fleming’s James Bond quartermaster Q. During World War II, Fraser-Smith worked for the Ministry of Supply, fabricating equipment nicknamed “Q-devices" (after Q-ships) for SOE agents operating in occupied Europe. Prior to the war, Fraser-Smith had worked as a missionary in North Africa. After the war he purchased a dairy farm in Bratton Fleming, Devon, where he died in 1992.
@@
LLM2GR: Ian Fleming created the fictional character of James Bond as the central figure for his works. Bond is an intelligence officer in the Secret Intelligence Service, commonly known as MI6. Bond is known by his code number, 007, and was a Royal Naval Reserve Commander. Fleming based his fictional creation on a number of individuals he came across during his time in the Naval Intelligence Division and 30 Assault Unit during the Second World War, admitting that Bond “was a compound of all the secret agents and commando types I met during the war". Among those types were his brother, Peter, who had been involved in behind-the-lines operations in Norway and Greece during the war.
Table 12: In the HotpotQA dataset, the Gold Standard, BM25, and LLM2GR reference passage for the query “Who is the fictional head of a British Secret Service division and for which a one-time missionary was the inspiration for?" are provided. The parts containing the answer are highlighted with a grey background.

### v3

@@ B.3 Further Analysis
Impact of Document Selection ($k$) of First Stage Documents. In Figure 7, we conduct experiments to compare the effect of selecting different numbers of first-stage documents (denoted as $k$) in MindRef when implemented based on LLaMA-2-13b. We observe that the impact of $k$ on the final performance is not significant, as the necessary effective reference passages are usually contained within the first few documents. The suboptimal performance observed on the HotpotQA dataset when $k=1$ can be attributed to the dataset requiring two documents to calculate R-Precision.
@@
Impact of Beam Search Sizes. Figures 8 and 9 show the impact of setting different beam sizes in the first and second stages, respectively, in MindRef when implemented based on LLaMA-2-13b. For the first stage of recalling title identifiers, a larger beam size can achieve better page-level results, thereby slightly improving the effectiveness of the second stage of passage recall. However, in the second stage of fine-grained passage recall, the improvement brought by a larger beam size is not significant and may even lead to a slight decline, possibly due to the introduction of additional noise by a larger beam size.
Memory Usage Analysis. Dense retrieval methods such as Contriever and DPR require over 60GB of memory usage. In contrast, sparse retrieval methods use far less memory, only needing 17GB. The MindRef framework, utilizing FM-index and Trie indexing, requires only 8GB when pre-encoding and storing all documents with FM-index, and the Trie storing all title identifiers needs just 25MB, which is negligible. Compared to sparse and dense retrieval methods, the recall framework effectively saves memory.
B.4 Case Study
@@
In Supplementary Material Tables 7 to 12, we present reference cases obtained using the Gold Standard, BM25, and the MindRef framework with LLaMA-2-13b on the NQ, TriviaQA, and HotpotQA datasets. By generating passages more aligned with the question, MindRef achieves results that contain the answer in Supplementary Material Tables 7, 9, and 11. Supplementary Material Table 8 showcases a biology question; although the passage recalled and located by MindRef does not contain the annotated answer, it provides a more detailed description of the location and process of pancreatic enzyme cleavage of peptide bonds. However, Supplementary Material Tables 10 and 12 show instances where MindRef’s recall failed. This is because merely generating a relevant prefix sometimes cannot ensure that the subsequent part will definitely contain the answer, leading to passages that are only broadly related. Ensuring the flexibility of recall and location while considering more subsequent information still requires further exploration. Nevertheless, we can also note that the references obtained by LLM recall are more natural and easier to read compared to those with predefined segmented beginnings. Finally, compared to the NQ dataset, questions in the TriviaQA and HotpotQA datasets are longer, allowing for more vocabulary overlap between questions and passages, which enhances the effectiveness of BM25 sparse retrieval.
@@
Gold Standard: Part II (s.10-21) deals with the marriageable age and the marriage of minors. In the original 1961 Act, marriageable age was set at 16 for females and 18 for males. However, under section 12 of the original 1961 Act a female 14 or 15 years or a male 16 or 17 years could apply to the court for permission to marry. The marriageable age was equalised in 1991 by the "Sex Discrimination Amendment Act 1991", which raised the marriageable age of females to 18.
BM25: marriage. Before 2004, there was no definition of marriage in the 1961 Act, and instead the common law definition used in the English case "Hyde v Hyde" (1866) was considered supreme. Though s.46(1) of the Act required celebrants to explain the legal nature of marriage in Australia to a couple as "the union of a man and a woman to the exclusion of all others, voluntarily entered into for life", these words were descriptive or explanatory, rather than outlining what constituted a legally valid marriage in
MindRef: Marriage in Australia is regulated by the federal "Marriage Act 1961" (Cth) , which applies uniformly throughout Australia (including its external territories) to the exclusion of all state laws on the subject. Australian law recognises only monogamous marriages, being marriages of two people, including same-sex marriages, and does not recognise any other forms of union, including traditional Aboriginal marriages, polygamous marriages or concubinage. The marriage age for marriage in Australia is 18 years, but in "unusual and exceptional circumstances" a person aged 16 or 17 can marry with parental consent
Table 7: In the NQ dataset, the Gold Standard, BM25, and MindRef reference passage for the query "what is the legal age for marriage in Australia" are provided. The parts containing the answer are highlighted with a grey background.
@@
Gold Standard: "In vivo", chymotrypsin is a proteolytic enzyme (serine protease) acting in the digestive systems of many organisms. It facilitates the cleavage of peptide bonds by a hydrolysis reaction, which despite being thermodynamically favorable, occurs extremely slowly in the absence of a catalyst. The main substrates of chymotrypsin are peptide bonds in which the amino acid N-terminal to the bond is a tryptophan, tyrosine, phenylalanine, or leucine. Like many proteases, chymotrypsin also hydrolyses amide bonds "in vitro", a virtue that enabled the use of substrate analogs such as N-acetyl-L-phenylalanine p-nitrophenyl amide for enzyme assays.
BM25: 149, producing $\alpha$-chymotrypsin (which is more active and stable than $\pi$-chymotrypsin). The resulting molecule is a three-polypeptide molecule interconnected via disulfide bonds. "In vivo", chymotrypsin is a proteolytic enzyme (serine protease) acting in the digestive systems of many organisms. It facilitates the cleavage of peptide bonds by a hydrolysis reaction, which despite being thermodynamically favorable, occurs extremely slowly in the absence of a catalyst. The main substrates of chymotrypsin are peptide bonds in which the amino acid N-terminal
MindRef: Chymotrypsin cleaves peptide bonds by attacking the unreactive carbonyl group with a powerful nucleophile, the serine 195 residue located in the active site of the enzyme, which briefly becomes covalently bonded to the substrate, forming an enzyme-substrate intermediate. Along with histidine 57 and aspartic acid 102, this serine residue constitutes the catalytic triad of the active site. These findings rely on inhibition assays and the study of the kinetics of cleavage of the aforementioned substrate, exploiting the fact that the en
Table 8: In the NQ dataset, the Gold Standard, BM25, and MindRef reference passage for the query "where does cleavage of the peptide bond by chymotrypsin occur" are provided. The parts containing the answer are highlighted with a grey background.
@@
Gold Standard: Only 193 of the estimated 643 people on board survived, and the soldiers’ chivalry gave rise to the unofficial "women and children first" protocol when abandoning ship, while the "Birkenhead drill" of Rudyard Kipling’s poem came to describe courage in face of hopeless circumstances.
BM25: HMS "Birkenhead, also referred to as HM Troopship "Birkenhead or Steam Frigate "Birkenhead", was one of the first iron-hulled ships built for the Royal Navy. She was designed as a steam frigate, but was converted to a troopship before being commissioned. She was wrecked on 26 February 1852, while transporting troops to Algoa Bay at Danger Point near Gansbaai, 87 miles (140 kilometres) from Cape Town in the Cape Colony. There were not enough serviceable lifeboats for all the passengers, and the soldiers
MindRef: The sinking of the "Birkenhead" is one of the earliest maritime disaster evacuations during which the concept of "women and children first" is known to have been applied."Women and children first" subsequently became standard procedure in relation to the evacuation of sinking ships, in fiction and in life. The term "Birkenhead drill" became defined as courageous behaviour in hopeless circumstances and appeared in Rudyard Kipling’s 1893 tribute to the Royal Marines, "Soldier an’ Sailor Too": To take your chance in the thick of a rush, with firing all about, Is nothing so bad when you’ve
Table 9: In the TriviaQA dataset, the Gold Standard, BM25, and MindRef reference passage for the query "Which expression is associated with the sinking of the HMS Birkenhead at Gansbaai near Cape Town, South Africa, in Febuary 1852?" are provided. The parts containing the answer are highlighted with a grey background.
@@
Gold Standard: The first series of I’m a Celebrity…Get Me Out of Here! was broadcast on ITV from 25 August to 8 September 2002. Ant & Dec presented the main show on ITV, whilst Louise Loughman hosted the spin-off show "I’m a Celebrity…Get Me Out of Here! NOW!" on ITV2. The winner of this series was radio DJ Tony Blackburn.
BM25: The first series of I’m a Celebrity…Get Me Out of Here! was broadcast on ITV from 25 August to 8 September 2002. Ant & Dec presented the main show on ITV, whilst Louise Loughman hosted the spin-off show "I’m a Celebrity…Get Me Out of Here! NOW!" on ITV2. The winner of this series was radio DJ Tony Blackburn. The show began with 8 celebrity contestants. The contestants take part in daily trials to earn food All ratings are taken from the UK
MindRef: I’m a Celebrity…Get Me Out of Here! is a British reality TV series in which a number of celebrities live together in a jungle environment for a number of weeks, competing to be crowned "King" or "Queen of the Jungle". The show was originally created in the United Kingdom by the factual programmes department of ITV’s then London franchise, London Weekend Television and developed by a team including James Allen, Natalka Znak, Brent Baker and Stewart Morris. The first episode aired on 25 August 2002 . It is now produced by ITV Studios and has been licensed globally to countries including the United States
Table 10: In the TriviaQA dataset, the Gold Standard, BM25, and MindRef reference passage for the query "Who was the first winner of ‘I’m A Celebrity, Get Me Out Of Here’?" are provided. The parts containing the answer are highlighted with a grey background.
@@
BM25: S is a South Korean project group consisting of three members: Kangta, Lee Ji-hoon and Shin Hye-sung. The group debuted in 2003, under the SM Entertainment label. After 11 years, they released and promoted another mini-album in 2014. In 2003, S released their first album "Fr.In.Cl", which stands for "Friends in Classic". In 2014, after more than a decade in hiatus, the group released their second mini-album "Autumn
MindRef: 2014 S/S is the debut album of South Korean group WINNER. It was released on August 12, 2014 by the group’s record label, YG Entertainment. The members were credited for writing the lyrics and composing the majority of the album’s songs. The members produced the majority of the music for this album themselves, with the help of other producers such as Choice 37, B.I., Airplay, and others. The album was highlighted for incorporating elements generally absent from K-pop releases, including hints of acoustic
Table 11: In the HotpotQA dataset, the Gold Standard, BM25, and MindRef reference passage for the query "2014 S/S is the debut album of a South Korean boy group that was formed by who?" are provided. The parts containing the answer are highlighted with a grey background.
@@
Gold Standard: Charles Fraser-Smith (26 January 1904 – 9 November 1992) was an author and one-time missionary who is widely credited as being the inspiration for Ian Fleming’s James Bond quartermaster Q. During World War II, Fraser-Smith worked for the Ministry of Supply, fabricating equipment nicknamed "Q-devices" (after Q-ships) for SOE agents operating in occupied Europe. Prior to the war, Fraser-Smith had worked as a missionary in North Africa. After the war he purchased a dairy farm in Bratton Fleming, Devon, where he died in 1992.
@@
MindRef: Ian Fleming created the fictional character of James Bond as the central figure for his works. Bond is an intelligence officer in the Secret Intelligence Service, commonly known as MI6. Bond is known by his code number, 007, and was a Royal Naval Reserve Commander. Fleming based his fictional creation on a number of individuals he came across during his time in the Naval Intelligence Division and 30 Assault Unit during the Second World War, admitting that Bond "was a compound of all the secret agents and commando types I met during the war". Among those types were his brother, Peter, who had been involved in behind-the-lines operations in Norway and Greece during the war.
Table 12: In the HotpotQA dataset, the Gold Standard, BM25, and MindRef reference passage for the query "Who is the fictional head of a British Secret Service division and for which a one-time missionary was the inspiration for?" are provided. The parts containing the answer are highlighted with a grey background.